# ConDiSim: Conditional Diffusion Models for Simulation-Based Inference

### v1

@@ 5.2 Scalability
ConDiSim’s scalability was extensively evaluated across diverse benchmark problems, spanning from low-dimensional (Two Moons, GMM, SIR; 2 parameters) to higher-dimensional scenarios (Gaussian Linear, Bernoulli GLM; 10 parameters, HH Model; 7 parameters, genetic oscillator; 15 parameters). From the C2ST metric values in Table 1, and the posterior plots spanning from Figure 2(a) to Figure 13, it can be seen that ConDiSim maintained the posterior approximation accuracy with increasing problem dimensionality and complexity. Model training times in Table 2 also show that ConDiSim is computationally efficient to train and scales favorably with problem complexity, making it suitable for high-dimensional SBI tasks. Its ability to handle larger parameter spaces without significant loss of performance is a key advantage, as evidenced in the genetic oscillator test problem involving 15 parameters.
5.3 Comparative Evaluation of different Inference Methods
Tables 1 and 4 indicate that diffusion models (Simformer, ConDiSim) generally outperform flow-based (NPE, APT) and GAN-based (GATSBI) approaches in accuracy, owing to their ability to model complex posteriors. Simformer achieves more accurate posterior approximations in most test problems (also see Table 4) owing to its complex architecture, however it also entails longer training times due to the involved complexity (Table 2). In contrast, ConDiSim’s simplified conditioning mechanism and architecture reduces training complexity, and its iterative denoising process does not impede inference speed in our experiments. The known challenges of adversarial training affect GATSBI, while APT’s sequential nature increases its computational cost relative to NPE.
@@ 5.4 Robustness
Robustness to observational noise is a critical factor in the practical application of simulation-based inference (SBI) methods. To rigorously evaluate this aspect, we conducted experiments on two challenging tasks: the SLCP Distractors and the Bernoulli GLM Raw. The SLCP Distractors task introduces 92 non-informative variables to the observations, significantly increasing the inference complexity. In contrast, the Bernoulli GLM Raw task utilizes raw data instead of well-behaved summary statistics, presenting a more realistic scenario often encountered in real-world applications. The results for the SLCP Distractors and Bernoulli GLM RAW, illustrated in Table 1 and Figures 3(b),12(b) demonstrate ConDiSim’s robustness to both observational noise. Such resilience is important as it indicates ConDiSim’s capacity to effectively handle real-world data, which is often characterized by measurement errors and inherent uncertainties.
@@ 5.5 Reliability
To further validate the learned inference model, we employed Simulation-Based Calibration (SBC) as a rigorous diagnostic for assessing the quality of approximate posterior distributions. The empirical cumulative distribution function (ECDF) plots demonstrate ConDiSim’s ability to produce accurate and well-calibrated posteriors, with rank statistics closely aligning along the diagonal—indicative of unbiased and reliable inference. This alignment supports the statistical validity and robustness of ConDiSim’s inferences across diverse and challenging scenarios.
@@ 6 Conclusion
In this work, we introduced ConDiSim, a denoising conditional diffusion model for simulation-based inference. We demonstrated ConDiSim’s ability to effectively capture complex posterior distributions efficiently on 10 benchmark problems and 2 real-world test problems. ConDiSim performs comparably to the state-of-the-art while being substantially faster to train. As part of an ablation study, we also explored the significance of the variance schedule and the number of diffusion steps in influencing model performance. ConDiSim offers a flexible and efficient approach to simulation-based inference, with potential applications across various scientific and engineering domains. Future work includes exploring adaptive variance schedules, integrating domain-specific knowledge into the model, and extending ConDiSim towards learning both the likelihood and posterior distributions.
@@ Acknowledgments
Appendix A SBI Problem Description
@@ A.1 Two Moons
  • Prior: $\bm{\theta}\sim\mathcal{U}(-1,1)$
@@
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$
A.2 Gaussian Mixture
@@
  • Prior: $\bm{\theta}\sim\mathcal{U}(-10,10)$
  • Simulator: $\mathbf{y}|\bm{\theta}=0.5\left(\mathcal{N}(\bm{\theta},\mathbf{I})+\mathcal{N}(\bm{\theta},0.01\mathbf{I})\right)$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$
A.3 Gaussian Linear
@@
  • Prior: $\bm{\theta}\sim\mathcal{N}(\mathbf{0},0.1\mathbf{I})$
  • Simulator: Generates data $\mathbf{y}$ using $\mathbf{y}|\bm{\theta}\sim\mathcal{N}(\bm{\theta},0.1\mathbf{I})$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$
A.4 Gaussian Linear Uniform
@@ A.5 SLCP
  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$
@@
    $y_{i}\sim\mathcal{N}\left(\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix},\begin{bmatrix}\theta_{3}^{2}&\tanh(\theta_{5})\theta_{3}\theta_{4}\\ \tanh(\theta_{5})\theta_{3}\theta_{4}&\theta_{4}^{2}\end{bmatrix}\right)$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{8}$
A.6 SLCP with Distractors
@@
  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$
@@
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{100}$
A.7 Bernoulli GLM
@@
  • Prior: $\beta\sim\mathcal{N}(\mathbf{0},2)$, $f\sim\mathcal{N}(\mathbf{0},(\mathbf{F}^{\top}\mathbf{F})^{-1})$
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$, with $\eta(z)=\frac{\exp(z)}{1+\exp(z)}$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$
  • Fixed Parameters: $T=100$
A.8 Bernoulli GLM Raw
@@
  • Prior: Same as Bernoulli GLM
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{100}$
  • Fixed Parameters: $T=100$
A.9 SIR Model
@@
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\mathcal{B}(1000,\frac{I}{N})$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{10}$
  • Fixed Parameters: Population size $N=1,000,000$, task duration $T=160$
A.10 Lotka-Volterra
@@
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\text{LogNormal}(\log(X),0.1)$ and $y_{2i}\sim\text{LogNormal}(\log(Y),0.1)$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{4},\mathbf{y}\in\mathbb{R}^{20}$
  • Fixed Parameters: Task duration $T=20$, initial conditions $(X(0),Y(0))=(30,1)$
A.11 Hodgkin-Huxley Model

### v2

@@ 5.2 Scalability
We evaluate ConDiSim across a spectrum of SBI benchmarks, ranging from low-dimensional problems (Two Moons, GMM, SIR; 2 parameters) to higher-dimensional settings (Gaussian Linear, Bernoulli GLM; 10 parameters, HH Model; 7 parameters, genetic oscillator; 15 parameters). As shown by C2ST scores in Table 1 and posterior visualizations (Figures 2(a)13), ConDiSim maintains high posterior fidelity even as dimensionality and model complexity increase. Training times in Table 2 further confirm its computational efficiency, with favorable scaling relative to problem size. Notably, the 15-dimensional genetic oscillator highlights ConDiSim’s architectural efficiency and effectiveness, as it preserves inference accuracy without suffering substantial degradation owing to a relatively large parameter space and the noisy, stochastic nature of the simulator.
@@ 5.3 Robustness
In many real-world problems, observational data are inherently imperfect and often contaminated by measurement noise, confounded by irrelevant features, or available only in raw, unprocessed form. Consequently, it is essential to evaluate how inference methods perform under such adverse conditions. We assess ConDiSim’s robustness on two challenging benchmarks: SLCP Distractors, where 92 uninformative variables are added to the observation vector to simulate high-dimensional nuisance features, and Bernoulli GLM Raw, which operates directly on raw binary observations rather than relying on handcrafted summary statistics, emulating realistic scenarios with noisy, unprocessed measurements. The genetic oscillator test problem is also noisy, simulated using the Stochastic Simulation Algorithm (SSA) [Gillespie, 2007]. As shown in Table 1 and Figures 3(b)12(b), ConDiSim maintains accurate posterior recovery across both settings, demonstrating resilience to observation noise and feature redundancy that demonstrates its suitability for practical scientific applications where data are high-dimensional, unprocessed, or corrupted by measurement uncertainty.
@@ 5.4 Reliability
Beyond predictive accuracy, it is crucial to verify that the learned posteriors are statistically well-calibrated, meaning that the reported uncertainty faithfully captures the true variability of the inference problem. To assess calibration, we use Simulation-Based Calibration (SBC) [Talts et al., 2018], which is a standard diagnostic that evaluates whether the rank statistics of posterior samples follow their expected uniform distribution. The empirical cumulative distribution function (ECDF) differences shown in Figures 2(b)15 remain within the expected bounds, with fluctuations centered around zero, indicating that ConDiSim produces unbiased and well-calibrated posteriors across benchmarks. Such calibration is essential for trustworthy inference in scientific applications, where both accurate point estimates and reliable uncertainty quantification are required.
@@ 6 Conclusion
We introduce ConDiSim, a DDPM-based conditional diffusion model for simulation-based inference. Across ten benchmark tasks and two real-world problems, ConDiSim effectively handles ill-posed inference, multimodal and high-dimensional posteriors, and demonstrates robustness even with raw, unprocessed observations, or in the presence of noise and distractors. It also adapts naturally to diverse data modalities, while offering substantially faster training time than competing methods. We demonstrate that ConDiSim is particularly suited for inference of stochastic simulators with time series output, where it outperforms other methods and delivers accuracy, well-calibrated posterior estimates, and fast amortized inference.
@@ Acknowledgments
Appendix A Connection between ConDiSim and Score based Diffusion Model
In this section, we establish a formal theoretical connection between ConDiSim, a DDPM based diffusion model, and the score based diffusion models. We demonstrate its equivalence to a variance preserving stochastic differential equation (SDE) by deriving the continuous time limit of ConDiSim’s discrete forward process, providing a unified perspective on diffusion based generative modeling.
A.1 Forward Process and Continuous-Time Limit
ConDiSim employs a discrete time forward process, defined as:
$\bm{\theta}_{t}=\sqrt{1-\beta_{t}}\,\bm{\theta}_{t-1}+\sqrt{\beta_{t}}\,\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$ (17)
where $\bm{\theta}_{t}$ represents the state at time step $t$, $\beta_{t}$ is the noise schedule, and $\bm{\epsilon}$ is standard Gaussian noise. To convert this discrete time process to continuous time, we consider the limit as the number of time steps $T\to\infty$ and define a small time step $\Delta t=\frac{1}{T}\to 0$. We introduce a continuous noise schedule $\beta(t)$ such that $\beta_{t}\approx\beta(t)\Delta t$. Using a first-order Taylor expansion for small $\beta_{t}$, we approximate $\sqrt{1-\beta_{t}}\approx 1-\frac{1}{2}\beta(t)\Delta t$. Substituting into the forward process:
$\bm{\theta}_{t}=\left(1-\frac{1}{2}\beta(t)\Delta t\right)\bm{\theta}_{t-1}+\sqrt{\beta(t)\Delta t}\,\bm{\epsilon},$ (18)
Defining the increment $\Delta\bm{\theta}_{t}=\bm{\theta}_{t}-\bm{\theta}_{t-1}$, and noting that the noise term $\sqrt{\beta(t)\Delta t}\,\bm{\epsilon}$ corresponds to a Wiener process increment $\Delta\mathbf{W}\sim\mathcal{N}(\mathbf{0},\Delta t\mathbf{I})$, we rewrite the equation as:
$\Delta\bm{\theta}_{t}=-\frac{1}{2}\beta(t)\bm{\theta}_{t-1}\Delta t+\sqrt{\beta(t)}\,\Delta\mathbf{W}.$ (19)
In the continuous-time limit ($\Delta t\to 0$), we replace $\Delta\bm{\theta}_{t}\to d\bm{\theta}_{t}$, $\Delta t\to dt$, $\Delta\mathbf{W}\to d\mathbf{w}_{t}$, and $\bm{\theta}_{t-1}\simeq\bm{\theta}_{t}$, yielding the following stochastic differential equation (SDE):
$d\bm{\theta}_{t}=-\frac{1}{2}\beta(t)\bm{\theta}_{t}\,dt+\sqrt{\beta(t)}\,d\mathbf{w}_{t},$ (20)
where $\mathbf{w}_{t}$ denotes standard Brownian motion and $d\mathbf{w}_{t}\sim\mathcal{N}(0,dt\,\mathbf{I})$. Comparing Equation 20 with the general form of the stochastic differential equations (SDEs):
$d\bm{\theta}_{t}=\mathbf{f}(\bm{\theta}_{t},t)\,dt+g(t)\,d\mathbf{w}_{t},$ (21)
where, $\mathbf{f}(\bm{\theta}_{t},t)$ is the deterministic drift term and $g(t)$ is the scalar valued diffusion coefficient. Equation (20) is a special case of the general SDE, with $\mathbf{f}(\bm{\theta}_{t},t)=-\frac{1}{2}\beta(t)\bm{\theta}_{t}$ and $g(t)=\sqrt{\beta(t)}$; this is known as the Variance-Preserving SDE (VP-SDE) [Song et al., 2021] and corresponds to an Ornstein–Uhlenbeck (OU) process. The solution to the VP-SDE can be derived using Itô calculus and is given by:
$\bm{\theta}_{t}=\sqrt{\alpha_{t}}\,\bm{\theta}_{0}+\sqrt{1-\alpha_{t}}\,\bm{\epsilon},\quad\text{where }\alpha_{t}=\exp\left(-\int_{0}^{t}\beta(s)\,ds\right),\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}).$ (22)
This expression represents the marginal distribution $q(\bm{\theta}_{t}\mid\bm{\theta}_{0})$ under the continuous diffusion process and serves as the continuous time analog of the discrete DDPM forward process:
$\bm{\theta}_{t}=\sqrt{\alpha_{t}}\,\bm{\theta}_{0}+\sqrt{1-\alpha_{t}}\,\bm{\epsilon},\quad\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i}).$ (23)
Thus, the VP-SDE generalizes the discrete Markov chain of ConDiSim to continuous time.
A.2 Reverse Diffusion Process and Score Matching
Given a general forward SDE as in Equation 21, the associated reverse-time SDE describes the dynamics when running the process backward in time (from $t=T$ to $t=0$) and corresponds to the reverse diffusion process of ConDiSim. Following [Anderson, 1982], the general form of the reverse SDE is:
$d\bm{\theta}_{t}=\left[\mathbf{f}(\bm{\theta}_{t},t)-g^{2}(t)\,\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t})\right]dt+g(t)\,d\mathbf{w}_{t},$ (24)
where $q_{t}(\bm{\theta}_{t})$ is the marginal probability density of $\bm{\theta}_{t}$ at time $t$ and $d\mathbf{w}_{t}$ is Brownian motion in reverse time. For the VP-SDE (Equation (20)), the reverse SDE becomes:
$d\bm{\theta}_{t}=\left[-\frac{1}{2}\beta(t)\,\bm{\theta}_{t}-\beta(t)\,\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t})\right]dt+\sqrt{\beta(t)}\,d\mathbf{w}_{t}.$ (25)
In this formulation, all terms are known except for the score function $s(\bm{\theta}_{t},t)=\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t})$, which is crucial for simulating the reverse process. Approximating this score is a standard optimization problem known as score matching, which is well-studied in probabilistic sampling, especially in the context of Langevin dynamics. However, since the true score depends on the intractable marginal $q_{t}(\bm{\theta}_{t})$, we instead introduce a parameterized approximation $s_{\phi}(\bm{\theta}_{t},t)$ and learn it by minimizing the Fisher divergence (explicit score matching loss):
$\phi^{*}=\arg\min_{\phi}\;\mathbb{E}_{t}\;\mathbb{E}_{q_{t}(\bm{\theta}_{t})}\left[\left\|s_{\phi}(\bm{\theta}_{t},t)-\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t})\right\|^{2}\right].$ (26)
To address the intractability of the true score, several methods have been developed, including implicit score matching [Hyvärinen and Dayan, 2005], sliced score matching[Song et al., 2020], and denoising score matching [Vincent, 2011]. To align the training objective with ConDiSim, we adopt denoising score matching (DSM), which leverages the tractable conditional score $\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t}\mid\bm{\theta}_{0})$. The corresponding objective is given by:
$\phi^{*}=\arg\min_{\phi}\;\mathbb{E}_{t}\;\mathbb{E}_{q_{0}(\bm{\theta}_{0})}\;\mathbb{E}_{q_{t}(\bm{\theta}_{t}|\bm{\theta}_{0})}\left[\left\|s_{\phi}(\bm{\theta}_{t},t)-\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t}|\bm{\theta}_{0})\right\|^{2}\right].$ (27)
Given the closed-form solution for the forward process in Equation (22), the score function $\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t}\mid\bm{\theta}_{0})$ becomes tractable and can be written as:
$\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t}|\bm{\theta}_{0})=-\frac{\bm{\theta}_{t}-\sqrt{\alpha(t)}\,\bm{\theta}_{0}}{1-\alpha(t)}.$ (28)
Reparameterizing $\bm{\theta}_{t}=\sqrt{\alpha_{t}}\,\bm{\theta}_{0}+\sqrt{1-\alpha_{t}}\,\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, the DSM objective becomes:
$\phi^{*}=\arg\min_{\phi}\;\mathbb{E}_{t}\;\mathbb{E}_{q_{0}(\bm{\theta}_{0})}\;\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|s_{\phi}(\bm{\theta}_{t},t)+\frac{1}{\sqrt{1-\alpha_{t}}}\,\bm{\epsilon}\right\|^{2}\right].$ (29)
This objective aligns closely with the ConDiSim training objective, where the score is approximated as:
$s_{\phi}(\bm{\theta}_{t},t)\approx-\frac{1}{\sqrt{1-\alpha_{t}}}\hat{\bm{\epsilon}}_{\phi}(\bm{\theta}_{t},t),$ (30)
with $\hat{\bm{\epsilon}}_{\phi}(\bm{\theta}_{t},t)$ being the predicted noise. This equivalence demonstrates that ConDiSim’s training objective is a rescaled form of the DSM objective, unifying the discrete DDPM framework with continuous-time score-based diffusion models. This analysis reveals that, although ConDiSim is formulated as a discrete DDPM, it is fundamentally equivalent to a score-based diffusion model in the continuous-time limit. The VP-SDE provides a principled framework for analyzing ConDiSim’s dynamics: the forward diffusion process can be understood through the Fokker–Planck equation, which characterizes convergence properties of the marginal distributions, while the reverse diffusion process connects naturally to annealed Langevin dynamics.
Appendix B SBI Problem Description
@@ B.1 Two Moons
  • Prior: $\bm{\theta}\sim\mathcal{U}(-1,1)$.
@@
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$.
B.2 Gaussian Mixture
@@
  • Prior: $\bm{\theta}\sim\mathcal{U}(-10,10)$.
  • Simulator: $\mathbf{y}|\bm{\theta}=0.5\left(\mathcal{N}(\bm{\theta},\mathbf{I})+\mathcal{N}(\bm{\theta},0.01\mathbf{I})\right)$.
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$.
B.3 Gaussian Linear
@@
  • Prior: $\bm{\theta}\sim\mathcal{N}(\mathbf{0},0.1\mathbf{I})$.
  • Simulator: Generates data $\mathbf{y}$ using $\mathbf{y}|\bm{\theta}\sim\mathcal{N}(\bm{\theta},0.1\mathbf{I})$.
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$.
B.4 Gaussian Linear Uniform
@@ B.5 SLCP
  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$.
@@
[…]
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{8}$.
B.6 SLCP with Distractors
@@
  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$.
@@
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{100}$.
B.7 Bernoulli GLM
@@
  • Prior: $\beta\sim\mathcal{N}(\mathbf{0},2)$, $f\sim\mathcal{N}(\mathbf{0},(\mathbf{F}^{\top}\mathbf{F})^{-1})$.
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$, with $\eta(z)=\frac{\exp(z)}{1+\exp(z)}$.
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$.
  • Fixed Parameters: $T=100$.
B.8 Bernoulli GLM Raw
@@
  • Prior: Same as Bernoulli GLM.
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$.
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{100}$.
  • Fixed Parameters: $T=100$.
B.9 SIR Model
@@
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\mathcal{B}(1000,\frac{I}{N})$.
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{10}$.
  • Fixed Parameters: Population size $N=1,000,000$, task duration $T=160$.
B.10 Lotka-Volterra
@@
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\text{LogNormal}(\log(X),0.1)$ and $y_{2i}\sim\text{LogNormal}(\log(Y),0.1)$.
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{4},\mathbf{y}\in\mathbb{R}^{20}$.
  • Fixed Parameters: Task duration $T=20$, initial conditions $(X(0),Y(0))=(30,1)$.
B.11 Hodgkin-Huxley Model